# Soret and Dufour effects in hot and dense QCD matter

### v1

@@ Abstract
The gradients act as invisible engines of transport, converting microscopic imbalances into macroscopic flows, and thus providing deep insights into the dynamics of physical systems. Thermal gradients do not merely drive the flow of heat, but they also set the microscopic constituents of the system into motion. In such scenarios, the constituents of the system not only transport energy but also diffuse collectively under the influence of these gradients. For the very first time, we present a first-principles investigation of the Soret and Dufour effects in hot and dense quantum chromodynamics (QCD) matter. We use the relativistic Boltzmann transport equation under the relaxation time approximation. By incorporating chemical potential and temperature gradients into the kinetic theory framework, we derive explicit expressions for the Dufour coefficient, which quantifies the heat flow due to concentration gradients, and the Soret coefficient, which describes the particle diffusion induced by thermal gradients. These coupled-transport phenomena are traditionally studied in multi-component classical systems at low energy scales. In this study, we follow quasiparticle models for the deconfined phase and the hadron resonance gas model for the confined hadronic phase in the context of heavy-ion collisions. This study provides novel insights into the thermo-diffusion and diffusion-thermo phenomena and opens avenues for incorporating such effects in hydrodynamic modeling and transport simulations of QCD matter.
@@ I Introduction
The coupled-transport phenomena offer a unique window into the microscopic dynamics that govern the macroscopic behavior of the system. Such interesting coupled phenomena are the Dufour and Soret effects [39, 40, 41, 42, 43]. The Dufour effect refers to the generation of a heat current due to a concentration gradient, also known as the diffusion-thermo effect. The Soret effect describes a phenomenon where a temperature gradient leads to the diffusion of the constituents of the medium, hence called as thermo-diffusion effect. These effects have numerous practical applications across various fields, including chemical and geophysical engineering, which have attracted significant attention from researchers over the years [44]. For instance, Ref. [45] reported notable findings on the simultaneous transport of heat and mass through species inter-diffusion processes under the direct influence of the Soret and Dufour effects, particularly in natural convection driven by evaporation and condensation phenomena. Similarly, Ref. [46] incorporated multiple-relaxation mechanisms within the lattice Boltzmann framework to investigate dual-diffusive natural convective flows, demonstrating that the combined influence of Soret and Dufour effects plays a critical role in governing these transport processes. In the Ref. [47], analytical expressions for the Soret and Dufour effects were derived, highlighting their significant contribution to coupled-transport phenomena in magnetized electrodynamics plasmas. Studying these phenomena in the domain of high-energy physics opens an exciting opportunity for understanding the coupled-transport dynamics of QCD matter. In the realm of ultra-relativistic heavy-ion collisions, the local temperature gradient naturally arises as the system expands and cools, while concentration gradients appear due to the system being slightly away from equilibrium. Incorporating Dufour and Soret effects into the study of QCD matter would provide a new dynamical phenomenon that would affect the evolution of the medium. The Dufour effect influences the heat current in the medium due to the gradient in the baryon chemical potential, which can essentially affect the cooling of the medium. Moreover, the Soret coefficient modifies the diffusion current due to the existing temperature gradient and can eventually affect the diffusion of the baryons in the medium, which can introduce an effect on experimental observables such as net proton number fluctuations and charge particle spectra. Thus, a study of these coupled-transport effects in the domain of heavy-ion collisions is important and could provide valuable insights into the interplay between heat and particle transport in QCD matter.
@@ II.2 Ideal Hadron Resonance Gas Model
with coefficients $\alpha_{1}$ and $\alpha_{2}$ that quantify the strength of the electric field and temperature gradient in pushing the system out of equilibrium.
@@
After comparing the tensors of the same notations, we get
@@
Figure 1: Temperature dependence of scaled coefficients of coupled-transport in the hadron resonance gas (HRG) and quark-gluon plasma (QGP) phases at different values of baryon chemical potential $\mu_{B}$. Shown are: (top left) diffusion coefficient ($D/T$), (top right) Soret coefficient ($S_{T}/T$), (bottom left) Dufour coefficient ($D_{F}/T^{2}$), and (bottom right) thermal conductivity ($\kappa/T^{2}$). HRG results are plotted with solid lines and QGP results with dashed lines. Each colour corresponds to a fixed $\mu_{B}$ value.

### v2

@@ Abstract
The gradients act as invisible engines of transport, converting microscopic imbalances into macroscopic flows, and thus providing deep insights into the dynamics of physical systems. Thermal gradients do not merely drive the flow of heat, but they also set the microscopic constituents of the system into motion. In such scenarios, the constituents of the system not only transport energy but also diffuse collectively under the influence of these gradients. For the very first time, we present a first-principles investigation of the Soret and Dufour effects in hot and dense quantum chromodynamics (QCD) matter. We use the relativistic Boltzmann transport equation under the relaxation time approximation. By incorporating baryon chemical potential and temperature gradients into the kinetic theory framework, we derive explicit expressions for the Dufour coefficient, which quantifies the heat flow due to concentration gradients, and the Soret coefficient, which describes the particle diffusion induced by thermal gradients. These coupled-transport phenomena are traditionally studied in multi-component classical systems at low energy scales. In this study, we follow quasiparticle models for the deconfined phase and the hadron resonance gas model for the confined hadronic phase in the context of heavy-ion collisions. This study provides novel insights into the thermo-diffusion and diffusion-thermo phenomena and opens avenues for incorporating such effects in hydrodynamic modeling and transport simulations of QCD matter.
@@ I Introduction
The coupled-transport phenomena offer a unique window into the microscopic dynamics that govern the macroscopic behavior of the system. Such interesting coupled phenomena are the Dufour and Soret effects [39, 40, 41, 42, 43]. The Dufour effect refers to the generation of a heat current due to a concentration gradient, also known as the diffusion-thermo effect. The Soret effect describes a phenomenon where a temperature gradient leads to the diffusion of the constituents of the medium, hence called as thermo-diffusion effect. These effects have numerous practical applications across various fields, including chemical and geophysical engineering, which have attracted significant attention from researchers over the years [44]. For instance, Ref. [45] reported notable findings on the simultaneous transport of heat and mass through species inter-diffusion processes under the direct influence of the Soret and Dufour effects, particularly in natural convection driven by evaporation and condensation phenomena. Similarly, Ref. [46] incorporated multiple-relaxation mechanisms within the lattice Boltzmann framework to investigate dual-diffusive natural convective flows, demonstrating that the combined influence of Soret and Dufour effects plays a critical role in governing these transport processes. In the Ref. [47], analytical expressions for the Soret and Dufour effects were derived, highlighting their significant contribution to coupled-transport phenomena in magnetized electrodynamics plasmas. Studying these phenomena in the domain of high-energy physics opens an exciting opportunity for understanding the coupled-transport dynamics of QCD matter. In the context of relativistic hydrodynamics, a thermo-diffusion contribution to the baryon diffusion current emerges naturally once the baryon sector is formulated within a causal second-order (Israel–Stewart) framework. In the first-order Navier–Stokes theory used, for example, in Ref. [48], the baryon current is driven by only a single thermodynamic force, which prevents the appearance of an independent response to a temperature gradient. Any temperature dependence is implicitly folded into a single transport coefficient. By contrast, second-order hydrodynamics, as mentioned in Ref. [49], provides a relaxation equation in which gradients of temperature and baryon density act as distinct thermodynamic forces with their own transport coefficients. This causal structure naturally allows for a thermo-diffusion (Soret-type) term that is dynamically generated rather than fixed by first-order constitutive relations. Moreover, the Ref. [50], explicitly identifies a baryon–heat cross conductivity that generates the coefficient of thermo-diffusion in a second-order framework. That study demonstrates that a temperature gradient can independently drive baryon diffusion when off-diagonal transport coefficients are included, in full agreement with Onsager reciprocity.
In the realm of ultra-relativistic heavy-ion collisions, the local temperature gradient naturally arises as the system expands and cools, while concentration gradients appear due to the system being slightly away from equilibrium. Incorporating Dufour and Soret effects into the study of QCD matter would provide a new dynamical phenomenon that would affect the evolution of the medium. The Dufour effect influences the heat current in the medium due to the gradient in the baryon chemical potential, which can essentially affect the cooling of the medium. Moreover, the Soret coefficient modifies the diffusion current due to the existing temperature gradient and can eventually affect the diffusion of the baryons in the medium, which can introduce an effect on experimental observables such as net proton number fluctuations and charge particle spectra. Thus, a study of these coupled-transport effects in the domain of heavy-ion collisions is important and could provide valuable insights into the interplay between heat and particle transport in QCD matter.
@@ II.2 Ideal Hadron Resonance Gas Model
with the unknown coefficients $\alpha_{1}$ and $\alpha_{2}$ that quantify the strength of the electric field and temperature gradient in pushing the system out of equilibrium.
@@
After comparing the coefficients on both the sides, we get
@@
Figure 1: Temperature dependence of scaled coefficients of coupled-transport in the hadron resonance gas (HRG) and quark-gluon plasma (QGP) phases at different values of baryon chemical potential $\mu_{B}$. Shown are scaled (top left) diffusion coefficient ($D/T$), (top right) Soret coefficient ($S_{T}/T$), (bottom left) Dufour coefficient ($D_{F}/T^{2}$), and (bottom right) thermal conductivity ($\kappa/T^{2}$). HRG results are plotted with solid lines and QGP results with dashed lines. Each colour corresponds to a fixed $\mu_{B}$ value.